# Analytics, have some humility: a statistical view of fourth-down decision making

### v4

@@ 3 Fourth-down decision making under uncertainty
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. The win probability models that form the crux of traditional fourth-down decision procedures today are machine learning models fit from historical data. Football analysts today typically use these models because they see a massive dataset – $600,825$ plays with $240,380$ first-down plays in our play-by-play dataset – and think this is enough data to fit a highly accurate model using machine learning. However, this is not true because the binary win/loss response variable is noisy and highly correlated: every game has only one winner. Formally, the binary response variable $y_{i}$ of the $i^{th}$ play indicates whether the team with possession won the game. The response values are not independent, as all plays from the same game share the same draw of the response column. Thus the effective sample size is somewhere between the number of plays ($600,825$) and the number of non-tied games ($4,101$) in our dataset. This is likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yardline, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context, and as such we expect win probability point estimates to have high variance.
3.1 Simulation study: random walk football
To illustrate just how difficult it is to accurately fit a statistical win probability model from noisy and highly correlated observational data, we conduct a simulation study. We create a simplified random walk version of football in which the true win probability at each game-state is known. Then, we see how well a statistical ${\mathsf{WP}}$ model recovers the true win probability. These models find the general ${\mathsf{WP}}$ trend (i.e., they they have low bias) but are subject to substantial uncertainty: to achieve adequate frequentist marginal coverage, bootstrapped ${\mathsf{WP}}$ confidence intervals need to be substantially wide. As real football is exponentially more complex, its confidence intervals should be even wider.
Rules of random walk football. Random walk football begins at midfield. Each play, the ball moves left or right by one yardline with equal probability. If the ball reaches the left (right) end of the field, team one (two) scores a touchdown, worth $+1$ ($-1$) point. The ball resets to midfield after each touchdown. After $N$ plays, the game ends. If the game is still tied after $N$ plays, a fair coin is flipped to determine the winner. We include the formal mathematical specification of the game in Appendix E. We also explicitly compute true win probability as a function of time, field position, and score differential using dynamic programming in Appendix E.
Simulation methodology. One simulated simplified football dataset consists of $G$ games, each with $N$ plays per game. We use $L=4$ yardlines so that the average number of first-down plays between each score is similar to that of a real football game. $M=25$ times we simulate a training dataset ${\mathscr{D}}_{\text{train}}$ and a testing dataset ${\mathscr{D}}_{\text{train}}$, each of which is a draw of simplified football history. Each of the $M$ simulated training and testing datasets has the form
${\mathscr{D}}_{\text{train}}=\{(n,X_{gn},S_{gn},y_{gn}):n=1,...,N\text{ and }g=1,...,G\}.$ (4)
and
${\mathscr{D}}_{\text{test}}=\{(n,X_{gn},S_{gn},{\mathsf{WP}}_{gn}):n=1,...,N\text{ and }g=1,...,G\},$ (5)
respectively. For each play of game $g$ we record the timestep $n$, the field position $X_{gn}$, the score differential $S_{gn}$, a binary variable $y_{gn}$ indicating whether the team with possession wins the game, and the true win probability ${\mathsf{WP}}_{gn}$. The response variable $y$ is clustered and highly correlated, as each play within the same game shares the same random draw of $y$. From the training set we fit a machine learning model to estimate win probability as a function of timestep $n$, field position $x$, and score differential $s$,
$\widehat{{\mathsf{WP}}}(n,x,s)={\mathsf{XGBoost}}({\mathscr{D}}_{\text{train}})(y|n,x,s).$ (6)
We use half of the games from the training set as a validation set to tune the ${\mathsf{XGBoost}}$ model. We also create various bootstrapped win probability confidence intervals, discussed below, denoted $CI(n,x,s)$.
For each simulation we compute the mean absolute error between true and estimated win probability on the test set,
${\mathsf{MAE}}=\frac{1}{G\cdot N}\sum_{(n,g)\in{\mathscr{D}}_{\text{test}}}|\widehat{{\mathsf{WP}}}(n,x_{gn},s_{gn})-{\mathsf{WP}}_{gn}|.$ (7)
Also, for each simulation and each confidence interval method, we compute the pointwise marginal coverage
$\text{coverage}=\frac{1}{G\cdot N}\sum_{(n,g)\in{\mathscr{D}}_{\text{test}}}\mathbb{I}\left({\mathsf{WP}}_{gn}\in CI(n,x_{gn},s_{gn})\right)$ (8)
and mean length
mean $\text{mean length}=\frac{1}{G\cdot N}\sum_{(n,g)\in{\mathscr{D}}_{\text{test}}}|CI(n,x_{gn},s_{gn})|.$ (9)
We report the ${\mathsf{MAE}}$, CI marginal coverage, and CI mean length averaged over the $M$ simulations and include plus/minus twice their standard errors.
Bootstrap confidence interval methodology. We compare the coverage and lengths of ${\mathsf{WP}}$ confidence intervals produced by the standard bootstrap, cluster bootstrap, and randomized cluster bootstrap, averaged over the $M$ simulations. In the standard bootstrap, which assumes each row (play) of the dataset is independently drawn, each of $B$ bootstrapped datasets are formed by re-sampling $N$ plays uniformly with replacement. In the cluster bootstrap, each of $B$ bootstrapped datasets are formed by re-sampling $G$ games uniformly with replacement, keeping each observed play within each re-sampled game. Finally, in the randomized cluster bootstrap, each of $B$ bootstrapped datasets are formed by re-sampling $G$ games uniformly with replacement, and within each game re-sampling $N$ plays uniformly with replacement.
Then, for each bootstrap method, we fit a ${\mathsf{WP}}$ model ${\mathsf{WP}}_{b}$ to each bootstrapped dataset $b=1,...,B$. The $90\%$ confidence interval for the ${\mathsf{WP}}$ estimate at game-state ${\mathbf{x}}$ is defined by the $5^{th}$ and $95^{th}$ quantiles of ${\mathsf{WP}}_{1}({\mathbf{x}}),...,{\mathsf{WP}}_{B}({\mathbf{x}})$. We use $B=100$ in this simulation study. Also, to avoid substantially low coverage near the extremes (${\mathsf{WP}}\approx 0$ or ${\mathsf{WP}}\approx 1$), we widen our confidence intervals when $\widehat{{\mathsf{WP}}}<0.025$ to have a lower bound of 0 and when $\widehat{{\mathsf{WP}}}>0.975$ to have an upper bound of 1.
Initial simulation results. In Table 1 we report the initial results of our simulation study. In the second column of Table 1, each simulated dataset consists of $G=4,101$ games and $N=53$ plays per game, which matches the number of games and the average number of first-down plays in our dataset of real football plays. Each game in each of these datasets consists of $K=53$ plays per game that share the same outcome. In the third column of Table 1, each simulated dataset consists of $G=4,101\cdot 53$ games and $N=53$ plays per game. Then, we remove all but $K=1$ play per game so that each timestep $n$ has exactly $4,101$ corresponding i.i.d. rows in the dataset. In other words, those datasets consist of $217,353$ plays with i.i.d. outcomes. For the simulation studies in each column, we report the mean absolute error of win probability point estimates. We also report the marginal coverage and mean length of $90\%$ confidence intervals produced by our three bootstrap methods. Alongside each value, which is an average over the $M$ simulations, we report plus/minus twice its standard error.
$G=4,101$,  $N=53$,  $K=53$ $G=4,101\cdot 53$,  $N=53$,  $K=1$
MAE bt ${\mathsf{WP}}$ and $\widehat{\mathsf{WP}}$ $0.0159\pm 0.0005$ $0.0155\pm 0.0004$
$90\%$ SB CI coverage $0.60\pm 0.029$ $0.65\pm 0.011$
$90\%$ CB CI coverage $0.74\pm 0.029$
$90\%$ RCB CI coverage $0.80\pm 0.028$
$90\%$ SB CI mean length $0.028\pm 0.001$ $0.029\pm 0.001$
$90\%$ CB CI mean length $0.047\pm 0.001$
$90\%$ RCB CI mean length $0.056\pm 0.001$
Table 1: Initial simulation study results for $G=4,101$ $N=53$, and $K=53$ (second column) and $G=4,101\cdot 53$ $N=53$, and $K=1$ (third column). We report the ${\mathsf{MAE}}$ (Formula (7)) of ${\mathsf{XGBoost}}$ ${\mathsf{WP}}$ point estimates and the marginal coverage (Formula (8)) and mean length (Formula (9)) of $90\%$ ${\mathsf{WP}}$ confidence intervals from the standard bootstrap (SB), cluster bootstrap (CB), and randomized cluster bootstrap (RCB). We include these values averaged over the $M$ simulations plus/minus twice their standard errors.
In the simulation study with highly correlated outcomes ($K=53$), the mean absolute error (${\mathsf{MAE}}$) between the true and estimated ${\mathsf{WP}}$ is less than $2\%$ over average. This indicates ${\mathsf{XGBoost}}$ recovers the general trend of true ${\mathsf{WP}}$, which we visualize in Figure 3a. In the simulation study with independent observations ($K=1$), the ${\mathsf{MAE}}$ is smaller. This suggests that some bias is induced by fitting ${\mathsf{WP}}$ from highly correlated observational data. Nonetheless, the ${\mathsf{MAE}}$ between the highly correlated ($K=53$) and independent ($K=1$) simulation studies are overall quite similar. This suggests that most of the error accrued when estimating ${\mathsf{WP}}$ from observational data is due to having limited data and a noisy binary response column.
Figure 3: Win probability ($y$-axis) as a function of play number $n$ ($x$-axis) for various values of score differential $s$ (color) at yardline $x=3$. True win probability is the solid line. On the left: for one of the simulations, ${\mathsf{WP}}$ estimates (dotted line) get the general trend right. On the right: $90\%$ bootstrapped ${\mathsf{WP}}$ confidence intervals (shaded regions) need to be substantially wide to achieve $90\%$ marginal frequentist coverage.
The marginal frequentist coverage of ${\mathsf{WP}}$ confidence intervals is also impacted by the correlated nature of football play-by-play data. The standard bootstrap achieves $5\%$ higher coverage for $K=1$ than for $K=53$. Since rows are independent in the simulation study without auto-correlation, each bootstrap method is identical when $K=1$. In the simulation study with highly correlated outcomes ($K=53$), mimicking observed football data, $90\%$ win probability confidence intervals yield much less than $90\%$ coverage. The best marginal coverage is a subpar $80\%$ from the randomized cluster bootstrap, whose $90\%$ confidence intervals have a large mean length $5.6\%$ ${\mathsf{WP}}$.
Tuning the bootstrap. Even in the obscenely simplified and idealized setting of random walk football, typical bootstrapped win probability confidence intervals don’t achieve proper coverage. The naive standard bootstrap in particular achieves dismally low coverage. Even the randomized cluster bootstrap that accounts for the correlation structure of observational football data doesn’t achieve high enough coverage. This coverage issue is likely worse for real football data, which is exponentially more complex than random walk football.
Since naive bootstrapped confidence intervals are not wide enough, we wish to tune the bootstrap so that its $90\%$ confidence intervals indeed produce $90\%$ marginal coverage. The traditional method of tuning non-parametric bootstrapped confidence intervals is to calibrate the bootstrapped quantiles (DiCiccio and Efron,, 1996). For instance, instead of using the $\alpha/2^{th}$ and $(1-\alpha/2)^{th}$ quantiles of ${\mathsf{WP}}_{1}({\mathbf{x}}),...,{\mathsf{WP}}_{B}({\mathbf{x}})$ to form a $1-\alpha$ confidence interval, we could use the $\beta/2^{th}$ and $(1-\beta/2)^{th}$ quantiles for some $\beta<\alpha$.
Although calibrating the bootstrapped quantiles fixes the coverage issue for ${\mathsf{WP}}$ confidence intervals, it doesn’t calibrate the bootstrap percentage, a metric that we introduce later in Section 3.2 to measure uncertainty in the estimated optimal fourth-down decision. Therefore, we use an alternate method to tune the bootstrap that calibrates both ${\mathsf{WP}}$ confidence intervals and the bootstrap percentage. Specifically, we introduce a tuning parameter $\phi\in(0,1]$ representing the fraction of data to be re-sampled in generating a bootstrapped dataset. By re-sampling less data than in the original dataset, we widen bootstrapped confidence intervals and increase marginal coverage. In the $\phi$-standard bootstrap, we re-sample $N\cdot\phi$ plays (rows) uniformly with replacement. In the $\phi$-cluster bootstrap, we re-sample $G\cdot\phi$ games uniformly with replacement, keeping each observed play within each re-sampled game. Finally, in the $\phi$-randomized cluster bootstrap, we re-sample $G\cdot\phi$ games uniformly with replacement, and within each game re-sample $N$ plays uniformly with replacement.
$\phi$-simulation results. In Table 2 we report the results of our simulation study in which we use $\phi$ to tune our bootstraps. Each simulated dataset consists of $G=4,101$ games, $N=53$ plays per game, and $K=53$ plays per game that share the same outcome. We report the marginal coverage and mean length of $90\%$ confidence intervals produced by our three bootstrap methods with $\phi=1$ (second column) and $\phi=0.5$ (third column). Alongside each value, which is an average over the $M$ simulations, we report plus/minus twice its standard error. As expected, lowering $\phi$ widens the confidence intervals and increases marginal coverage. As desired, $90\%$ confidence intervals from the randomized cluster bootstrap with $\phi=0.5$ achieve $90\%$ marginal coverage. To achieve satisfactory marginal coverage, however, ${\mathsf{WP}}$ confidence intervals need to be substantially wide, which we visualize in Figure 3b. Specifically, the randomized cluster bootstrap with $\phi=0.5$ produces $90\%$ confidence intervals with a mean width of $7.4\%$ ${\mathsf{WP}}$. This result is striking: in the extremely simplified setting of random walk football, win probability confidence intervals need to be extremely wide to achieve adequate coverage. So, it is very difficult to accurately estimate win probability by fitting a machine learning model from noisy and highly correlated football game outcomes. These are high variance estimators subject to large uncertainty.
$\phi=1$ $\phi=0.5$
$90\%$ $\phi$-SB CI coverage $0.60\pm 0.029$ $0.69\pm 0.027$
$90\%$ $\phi$-CB CI coverage $0.74\pm 0.029$ $0.85\pm 0.027$
$90\%$ $\phi$-RCB CI coverage $0.80\pm 0.028$ ${\color[rgb]{1,0,1}\textbf{0.90}}\pm 0.026$
$90\%$ $\phi$-SB CI mean length $0.028\pm 0.001$ $0.036\pm 0.001$
$90\%$ $\phi$-CB CI mean length $0.047\pm 0.001$ $0.063\pm 0.001$
$90\%$ $\phi$-RCB CI mean length $0.056\pm 0.001$ ${\color[rgb]{1,0,1}\textbf{0.074}}\pm 0.001$
Table 2: Simulation study results for $\phi=1$ (second column) and $\phi=0.5$ (third column) with $G=4,101$, $N=53$, and $K=53$ for both. We report the marginal coverage (Formula (8)) and mean length (Formula (9)) of $90\%$ ${\mathsf{WP}}$ confidence intervals from the standard bootstrap (SB), cluster bootstrap (CB), and randomized cluster bootstrap (RCB). We include these values averaged over the $M$ simulations plus/minus twice their standard errors.
3.2 Quantifying uncertainty in the estimated optimal decision
The fourth-down decision procedure widely used today in the football analytics community is to make the decision that maximizes estimated win probability. Win probability estimates arise from machine learning models fit from observational football data, which consists of noisy and highly correlated game outcomes. As discussed in the previous Section 3.1, even in simplified random walk football, win probability models are high variance estimators subject to substantial uncertainty. ${\mathsf{WP}}$ models fit from real football data are likely subject to even more uncertainty since real football is so much more complex. Therefore, it is imperative to incorporate uncertainty quantification into the fourth-down decision procedure.
Since we estimate win probability using machine learning models, we need a non-parametric method to quantify uncertainty in these estimates. Bootstrapping is a natural choice to capture such uncertainty. The bootstrapping process begins with generating $B$ bootstrapped datasets from our original dataset. We specify the full technical details of how we generate each bootstrapped dataset in Appendix F and how we choose an appropriate value for $B$ in Appendix G. Each bootstrapped dataset mimics a new draw of the training set: keeping all of the pre-game attributes of each game fixed (e.g., each team’s composition), bootstrapping mimics re-drawing plays and the random win/loss outcome of each game. Concisely, each bootstrapped dataset represents a re-draw of the recent history of football. Then, to each bootstrapped dataset we fit a fourth-down win probability model (described in Section 2.3 and Appendix B). At each game-state, each model yields an estimated optimal decision. The strength of each estimated optimal decision is proportional to the effect size, or the estimated gain in win probability by making that decision. Because the plays and win/loss outcomes in each bootstrapped dataset differ from those in the original dataset, each bootstrapped decision function differs from the original one.
Given the original effect size $\hat{g}({\mathbf{x}})$ and the $B$ bootstrapped effect sizes $\hat{g}_{1}({\mathbf{x}}),...,\hat{g}_{B}({\mathbf{x}})$ at game-state ${\mathbf{x}}$, we construct a confidence interval on the effect size using the quantiles. For instance, if $B=100$, a $90\%$ confidence interval on the effect size is $[\hat{g}_{5}({\mathbf{x}}),\ \hat{g}_{95}({\mathbf{x}})]$. This confidence interval represents the following. Assume that the win/loss outcomes across football history are generated by some underlying (and unknown) win probability function. If we re-simulated the history of football say $100$ times from that underlying ${\mathsf{WP}}$ function, keeping the pre-game conditions of each game the same, the true win probability is expected to lie in the $90\%$ confidence interval $90$ times.
Since a confidence interval is defined by two numbers, it is difficult to meld those two numbers into one decision recommendation. For example, consider a play in which the estimated gain in win probability for ${\mathsf{Go}}$ is $1\%$. If our confidence interval for this gain is $[-3\%,5\%]$, it is not clear how strongly we should recommend ${\mathsf{Go}}$. Similarly, it is difficult to compare two confidence intervals of varying lengths. For example, it is not clear whether our recommendation for a confidence interval of $[-3\%,5\%]$ should be weaker than that for a confidence interval of $[-1\%,3\%]$. Further, our primary interest is to quantify uncertainty in the decision itself (Friedman et al.,, 1999), which is not as granular as quantifying uncertainty in the effect size.
@@ 3.2 Quantifying uncertainty in the estimated optimal decision
To boil a bootstrapped confidence interval on the effect size into a single number to drive decision making, we bag the decision itself. Specifically, we consider the bootstrap percentage (${\mathsf{boot\%}}$), or the percentage of bootstrapped models that report each decision to be optimal. Formally, at game-state ${\mathbf{x}}$ we have the original estimated optimal decision $d({\mathbf{x}})\in\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ (the point estimate) and $B$ bootstrapped estimated optimal decisions $d_{1}({\mathbf{x}}),...,d_{B}({\mathbf{x}})$. Then
@@
Bootstrap percentage quantifies how confident we are that a fourth-down recommendation has higher win probability than all the other decisions. A high bootstrap percentage (${\mathsf{boot\%}}\approx 100\%$) reflects high confidence in the estimated optimal decision: the estimated optimal decision remains the same across the vast majority of re-draws of the training set. When ${\mathsf{boot\%}}$ is high, we can trust the decision recommended by the point estimate. Conversely, a low bootstrap percentage reflects low confidence in the estimated optimal decision: the estimated optimal decision differs substantially across different re-draws of the training set. A low bootstrap percentage is ${\mathsf{boot\%}}\approx 50\%$ or lower because a fourth-down decision almost always boils down to picking between ${\mathsf{Go}}$ and one of the kicks, as it is clearly sub-optimal to punt near the opponent’s endzone and to attempt a field goal far from the opponent’s endzone. Regardless of the effect size, when ${\mathsf{boot\%}}$ is low we cannot trust the point estimate: the effect size at that game-state is too dependent on the random idiosyncrasies of its particular training set. In other words, the edge detected by the point estimate is more due to noise than signal. In a “parallel universe” in which we re-drew each random win/loss game outcome, the estimated optimal decision at that game-state may differ solely due to the randomness inherent in drawing the history of football.
On this view, bootstrap percentage is a measure of data reliability. At each game-state the model produces a point estimate of the effectiveness of a fourth-down decision; bootstrap percentage tells us how reliable this estimate is, or how much the data trusts its own estimate. To understand, think of the outcome (winning team) of each row (play) in the dataset as a random draw. If some of these draws resulted in different outcomes, our fitted win probability functions would be different. The less data we have access to, the more sensitive models are to the random idiosyncrasies of any particular training dataset. The bootstrap quantifies this sensitivity: given the amount of data we have, it quantifies the spectrum of variability in potential resulting fitted models based only on sampling uncertainty, or uncertainty resulting from having limited data.

### v5

@@ 3 Fourth-down decision making under uncertainty
The standard approach, recommending the decision that maximizes estimated win probability, disregards the uncertainty that arises from estimating win probability from a finite dataset. Thus, in this section, we modify the fourth-down decision procedure to account for uncertainty.
Analysts may be inclined to trust estimates from win probability models because they are fit from an ostensibly massive dataset: there are $600,825$ plays and $229,635$ first-down plays in our play-by-play dataset. Nonetheless, several characteristics of this dataset reveal just how difficult it is to estimate win probability. First, the binary win/loss outcome variable is coarse. Since win probability is a continuous quantity in the interval $[0,1]$, to capture it with fine granularity from raw zeros and ones requires a massive amount of data. Estimating win probability as a function of game-state, with such a large set of possible game-states, dramatically increases the difficulty of the task. Win probability consists of complex nonlinear and interacting relationships between game-state variables, not simple additive relationships, which makes it even more difficult.
Moreover, the binary win/loss response variable is clustered: all plays from the same game share the same win/loss outcome. This dependence structure reduces the effective sample size, which is somewhere between the number of first-down plays ($229,635$) and the number of non-tied games ($4,101$) in our dataset. According to the simulation study in Brill and Wyner, (2024), we effectively have half as much data as we think. In other words, a win probability estimator fit from a dataset consisting of half as many plays, but where each play has an independent outcome, has the same accuracy as an estimator fit from a clustered dataset like our historical play-by-play dataset.
Thus, there is likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yardline, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context. Therefore, it is imperative to incorporate uncertainty quantification into the fourth-down decision procedure.
Since we estimate win probability using machine learning models, we need a non-parametric method to quantify uncertainty in these estimates. Bootstrapping is a natural choice to capture such uncertainty. The bootstrapping process begins with generating $B$ bootstrapped datasets from our original dataset. To do so, we use a randomized cluster bootstrap: re-sample games uniformly with replacement and within each game re-sample drives uniformly with replacement. According to the simulation study in Brill and Wyner, (2024), although the randomized cluster bootstrap achieves higher coverage than the standard i.i.d. bootstrap (re-sample plays uniformly with replacement), it produces undercovered intervals that are too narrow. Thus, we also tune our randomized cluster bootstrap using a fractional bootstrap as recommended in Brill and Wyner, (2024). We choose an appropriate value for $B$ ($B=101$) in Appendix C.
@@
Then, to each bootstrapped dataset, which represents a re-draw of the recent history of football, we fit a full fourth-down decision model (described in Section 2). At each game-state, each model yields an estimated optimal decision. To quantify uncertainty in the estimated optimal decision, we bag the decision itself. Specifically, we consider the bootstrap percentage (${\mathsf{boot\%}}$), or the percentage of bootstrapped models that report each decision to be optimal. Formally, at game-state ${\mathbf{x}}$ we have the original estimated optimal decision $d({\mathbf{x}})\in\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ (the point estimate) and $B$ bootstrapped estimated optimal decisions $d_{1}({\mathbf{x}}),...,d_{B}({\mathbf{x}})$. Then
@@
Bootstrap percentage quantifies how confident we are that a recommendation has higher win probability than the other decisions. A high ${\mathsf{boot\%}}$ ($\approx 100\%$) reflects high confidence: the estimated optimal decision remains the same across the vast majority of re-draws of the training set. Conversely, a low ${\mathsf{boot\%}}$ ($\approx 50\%$) reflects low confidence: the estimated optimal decision differs substantially across different re-draws of the training set. A low ${\mathsf{boot\%}}$ is $\approx 50\%$ or lower because a fourth-down decision almost always boils down to picking between ${\mathsf{Go}}$ and one of the kicks, as it is clearly sub-optimal to punt near the opponent’s endzone and to attempt a field goal far from the opponent’s endzone.
Regardless of the effect size, when ${\mathsf{boot\%}}$ is low we cannot trust the point estimate: the effect size at that game-state is too dependent on the random idiosyncrasies of its particular training set. In other words, the edge detected by the point estimate is more due to noise than signal. In a “parallel universe” in which we re-drew each random win/loss game outcome, the estimated optimal decision at that game-state may differ solely due to the randomness inherent in drawing the history of football.
We also consider confidence intervals on the effect size $g({\mathbf{x}})$, the gain in win probability by making a fourth-down decision at game-state ${\mathbf{x}}$. Given the original estimated effect size $\hat{g}({\mathbf{x}})$ and the $B$ bootstrapped effect sizes $\hat{g}_{1}({\mathbf{x}}),...,\hat{g}_{B}({\mathbf{x}})$, sorted from least to greatest, we construct a confidence interval on the effect size using the quantiles. For $B=101$, a $90\%$ confidence interval on the effect size is $[\hat{g}_{6}({\mathbf{x}}),\ \hat{g}_{96}({\mathbf{x}})]$. This confidence interval represents the following: assume that the win/loss outcomes across football history are generated by some underlying win probability function. If we re-simulated the history of football $B$ times from that ${\mathsf{WP}}$ function, keeping the pre-game conditions of each game the same, the true win probability is expected to lie in the $90\%$ confidence interval $0.9\cdot B$ times.
We find it easier for bootstrap percentage to drive decision making than confidence intervals because the former is defined by one number whereas the latter is defined by two numbers. For example, consider a play in which the estimated gain in win probability for ${\mathsf{Go}}$ is $1\%$. If our confidence interval for this gain is $[-3\%,5\%]$, it is not clear how strongly we should recommend ${\mathsf{Go}}$. Similarly, it is difficult to compare two confidence intervals of varying lengths. For example, it is not clear whether our recommendation for a confidence interval of $[-3\%,5\%]$ should be weaker than that for a confidence interval of $[-1\%,3\%]$. Also, our primary interest is to quantify uncertainty in the decision itself (Friedman et al.,, 1999), which is not as granular as quantifying uncertainty in the effect size.